star A is 4 magnitudes brighter than star B .how does the apparent brightness of star A compare with that of star B?


1
Expert's answer
2021-01-04T14:38:34-0500

According to the Pogson's ratio (see https://en.wikipedia.org/wiki/Apparent_magnitude#Calculations) the ratio of the apparent brightness of the stars is equal to:


"\dfrac{F_A}{F_B} \approx 2.512^{m_A-m_B}"

where "F_A", and "F_B" are apparent brightnesses of the stars A and B respectively, "m_A, m_B" are magnitudes of these stars. According to the problem statement "m_A-m_B = 4". Thus:


"\dfrac{F_A}{F_B} \approx 2.512^4\approx 39.8"

Answer. Star A is 39.8 times brighter than star B.